1.  A motorcyclist traveling with a constant speed of 15m/s passes a school crossing corner, where the speed limit is 10m/s. Just as the motorist passes the school crossing sign, a police officer on a motorcycle stopped there, starts in pursuit with a constant acceleration of 3.0m/s2.

I.  How much time elapses before the officer meets the motorist? 

II. What is the Police officer’s speed at that time? 

III. At that time, what distance have each vehicle traveled






1
Expert's answer
2021-03-23T08:13:13-0400

1)


"15t=0.5(3)t^2\\t=10\ s"


2)


"v=3(10)=30\frac{m}{s}"

3)


"S=15(10)=150\ m"